solve the following for x
2^x=32
3^x-2=81
4^x=1/64

Answer by stanbon(48535)

2^x=32
2^x = 2^5
x = 5
-------------
3^x-2=81
3^(x-2)=3^4
x-2 = 4
x = 6
--------------
4^x=1/64
4^x = 4^-3
x = -3
====================
Cheers,
Stan H.